OpenStudy (anonymous):

HELP FAST I WILL FAN AND MEDAL Find the measures of the complement and the supplement of an angle with a measure of 60°. A. complement, 30°; supplement, 120° B. complement, 330°; supplement, 120° C. complement, 160°; supplement, 160° D. complement, 220°; supplement, 70°

hartnn (hartnn):

compliment of an angle 'x' = 90 - x suppliment of an angle 'x' = 180 - x

hartnn (hartnn):

here x = 60 compliment = 90 - 60 = ... supplement = 180 - 60 = ...

OpenStudy (xguardians):

A

OpenStudy (anonymous):

Don't give out the answer , :(

OpenStudy (anonymous):

90-60= 30 180-60=120

hartnn (hartnn):

\(\huge \color{red}{\checkmark \quad \ddot \smile}\)